OpenStudy (anonymous):

Ah more physics help! A reconnaissance plane flies 531 km away from its base at 944 m/s, then flies back to its base at 1416 m/s. What is its average speed? Answer in units of m/s pretty please:)

OpenStudy (anonymous):

Average speed= total distance covered / elapsed time . This is an easy calculation.

OpenStudy (anonymous):

soo... 531000/ ? I don't understand how to get the elapsed time. Somehow i keep getting 566.4

OpenStudy (anonymous):

So you find the total time of the journey by v=d/t (be careful with your units make sure everything is in metres and seconds) and similarly for the flight back. You add those two times which give you total time. Total distance is just 531km(2)=1062km.

OpenStudy (anonymous):

Is it going to be 1062/(1416+944)? sorry i'm super bad at math and science

OpenStudy (anonymous):

Let's do this step by step. What's the time for the first trip?

OpenStudy (anonymous):

Well if it is 944m/s, is it going to be 944 m? Or is that supposed to be 531000/944?

OpenStudy (anonymous):

*s

OpenStudy (anonymous):

Yes so you get 562.5 seconds. What's the time for the second trip?

OpenStudy (anonymous):

375s?

OpenStudy (anonymous):

Correct. Now you have v= 2(531000m) / (375+562.5)s

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem :) .

OpenStudy (anonymous):

Do you mind helping me with one more problem?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

ok thanks. Consider a car which is traveling along a straight road with constant acceleration a .There are two checkpoints A and B which are a distance 108 m apart. The time it takes for the car to travel from A to B is 4.07 s. What is the velocity for B in the case of the acceleration being 3.82 m/s^2

OpenStudy (anonymous):

This is a 1-Dimensional problem . We are given a distance x of 108 m which the car covers in 4.07 seconds. We first find the initial speed at A from x=v(initial)t + 1/2at^2 . Then solve for v(final) from v(final)^2=v(initial)^2 + 2ax

OpenStudy (anonymous):

Ok, just to make sure I'm doing this right, is the initial 20.2 m/s?

OpenStudy (anonymous):

or 39.2?

OpenStudy (anonymous):

Don't keep guessing , derive the formula for v(initial). What is it?

OpenStudy (anonymous):

25.50

OpenStudy (anonymous):

What is v(initial)= ? in general...?

OpenStudy (anonymous):

You mean by definition? Starting velocity/ original velocity- in this case A is initial velocity

OpenStudy (anonymous):

Sorry I'm an idiot. But I did the formula of 108=4.07x + .5(3.08)(4.07)^2

OpenStudy (anonymous):

ok. I see what I did wrong. thank you for all your help!

OpenStudy (anonymous):

Don't mention it :)